Liquid hexane CH3(CH2)4CH3 will react with gaseous oxygen O2 to produce gaseous carbon dioxide CO2 and gaseous water H2O. Suppose 54.3 g of hexane is mixed with 73. g of oxygen. Calculate the minimum mass of hexane that could be left over by the chemical reaction. Be sure your answer has the correct number of significant digits. 


1
Expert's answer
2021-11-01T06:32:25-0400

The balanced reaction for the reaction of hexane with oxygen

"CH_3-(CH_2)4-CH_3 \ +9.5O_2\to 6CO_2+7H_2O"


Moles of hexane = 29/86=0.34

Moles of oxygen = 41.3/32=1.30

As we can see from the reaction that one mole of hexane requires 9.5 moles of oxygen

so oxygen is the limiting reagent here

moles of hexane consumed is 1.30/9=0.144

moles of hexane left = 0.34-0.14 =0.2 moles

mass of hexane left = 0.2×86=17.2 grams